### v1

# A continuum mechanics approach for the deformation of non-Euclidean origami generated by piecewise constant nematic director fields

Abstract
We merge classical origami concepts with active actuation by designing origami patterns whose panels undergo prescribed metric changes. These metric changes render the system non-Euclidean, inducing non-zero Gaussian curvature at the vertices after actuation. Such patterns can be realized by programming piecewise constant director fields in liquid crystal elastomer (LCE) sheets. In this work, we address the geometric design of both compatible reference director patterns and their corresponding actuated configurations. On the reference configuration, we systematically construct director patterns that satisfy metric compatibility across interfaces. On the actuated configuration, we develop a continuum mechanics framework to analyze the kinematics of non-Euclidean origami. In particular, we fully characterize the deformation spaces of three-fold and four-fold vertices and establish analytical relationships between their deformations and the director patterns. Building on these kinematic insights, we propose two rational designs of large director patterns: one based on a quadrilateral tiling with alternating positive and negative actuated Gaussian curvature, and another combining three-fold and four-fold vertices governed by a folding angle theorem. Remarkably, both designs achieve compatibility in both the reference and actuated states. We anticipate that our geometric framework will facilitate the design of non-Euclidean/active origami structures and broaden their application in active metamaterials, soft actuators, and robotic systems.
1 Introduction
Shape programming has emerged as a fundamental topic across diverse fields, including robotics, metamaterials, soft electronics, and geometric design choi2019programming ; filipov2015origami . Among various shape programming techniques, origami—the ancient art of paper folding—offers a promising platform for designing complex configurations through coordinated folding lines Levi2016Programming ; lang1996computational ; lang2011origami . Since paper is nearly inextensible, traditional origami involves isometric deformations that preserve distances within each panel. As a result, the surfaces remain developable, with zero Gaussian curvature, and the design space becomes highly constrained. The resulting nonlinearity and over-constrained geometry make many origami-related problems computationally challenging; for instance, determining rigid-foldability with optional creases is known to be NP-hard Akitaya_Demaine_Horiyama_Hull_Ku_Tachi_2020 . Due to these challenges, much research focuses on special classes of origami, such as rigidly and flat-foldable quadrilateral mesh origami feng2020designs , where inverse design frameworks have been developed dang2022inverse . Additional work on rigid foldability includes tachi2009generalization ; lang2018rigidly ; izmestiev2017classification ; hayakawa2024panel , offering accurate predictions for actuation, though typically requiring external control mechanisms ai2024easy .
An alternative approach to shape programming is metric design using responsive or functional materials. Systems inspired by biological growth huang , swelling gels HONG20093282 , or phase-transforming liquid crystal elastomers (LCEs) warner2007liquid can induce programmable metric changes (i.e., length changes) across a 2D domain. These metric changes directly influence the intrinsic geometry—such as Gaussian curvature—of the actuated surface klein ; mostajeran2015curvature ; mostajeran2016encoding . Several inverse design frameworks have emerged, including patterning LCEs aharoni2014geometry ; warner2018nematic ; itay2019curved ; griniasty2021multivalued , designing active nets aharoni2018universal , and controlling shape via pneumatic actuation siefert2019bio . The metric-induced curvature can also enhance the mechanical strength of the structure guin2018layered . These materials offer additional benefits such as remote control via light or magnetic fields kim2018printing , and their softness and biocompatibility make them suitable for integration in biomedical systems sekitani2016ultraflexible . However, the main limitation of the metric design approach is that it prescribes only the in-plane geometry (the first fundamental form); the out-of-plane shape (second fundamental form) must still be determined via minimization of a bending energy functional. This often results in flexible systems with highly implicit, energetically-selected deformations.
Figure 1: (a) An active origami adapted from plucinsky2018patterning . (b) A generalized design composed of three-fold and four-fold vertices with no symmetry.
Combining origami geometry with metric actuation seems to provide a new paradigm, potentially enabling shape-morphing systems that are both predictable, functionally programmable and remotely controllable. A seminal example is provided by Plucinsky et al. plucinsky2016programming ; plucinsky2018patterning (also shown in Fig. 1), where origami panels are augmented with panels patterned by piecewise-constant LCE director fields. Upon actuation, the panels elongate along the director by a factor $\lambda_{\parallel}$ and contract transversely by $\lambda_{\perp}$. Although each panel remains flat, the actuation induces non-isometric deformations, giving rise to non-trivial Gaussian curvature at the vertices o2006elementary ; duffy2020defective . The resulting configuration often bifurcates into distinct “up” and “down” states due to the induced curvature. While the experiments validate theoretical predictions, their design space remains limited due to symmetry constraints, limiting applicability in scenarios requiring more complex actuation. We also find that the systematic study of active non-Euclidean origami is quite limited in the literature. To enable more versatile functionality, it is crucial to develop a generalized design framework for active origami (also generically resulting in non-Euclidean origami waitukaitis2019non ) that can systematically handle director pattern design and accurately predict deformation behavior.
In this paper, we focus on the geometric and kinematic aspects of such a framework. We restrict our attention to a purely geometric regime, assuming that each panel undergoes a uniform, piecewise-constant metric change prescribed by the director field, and remains flat and rigid after actuation. Within this framework, two main challenges arise. First, the reference director pattern must be compatible, meaning the prescribed length changes must match across each crease. A simple case occurs when the crease is the bisector of adjacent directors or orthogonal to it—a condition referred to as metric compatibility feng2020evolving . However, patterning compatible director fields becomes nontrivial when accounting for topological and geometric interactions among multiple creases. Second, the kinematics of the actuated configuration must be predicted accurately. Ideally, the kinematics of the actuated shape can be predicted directly from the reference director pattern. A key question remains: even when the reference director pattern is compatible, does the actuated configuration maintain compatibility?
We address these challenges by first analyzing the fundamental conditions for metric compatibility and then studying the kinematics of actuated three-fold and four-fold intersections. We generalize previous continuum mechanics approaches feng2020designs ; zou2024kinematics to derive closed-form kinematic equations for these non-Euclidean vertices. Specifically, we show that the actuated three-fold vertices are rigid with no free folding angle and are fully determined by the director configuration, while the four-fold vertices have one free folding angle, with the remaining angles following from a unified analytical formula. Using these insights, we investigate multi-node interactions. For example, a pattern with two connected three-fold vertices reveals hidden degrees of freedom in director programming. For larger-scale patterns, we propose a marching algorithm to design quadrilateral director tilings with alternating positive and negative actuated Gaussian curvature. Additionally, we construct another class of compatible patterns composed of both three-fold and four-fold vertices (Fig. 1(b)), governed by a folding angle theorem. Surprisingly, despite the complexity of these systems, the active origami patterns are compatible in both the reference and actuated configurations and offer substantial design flexibility.
This paper is organized as follows. Section 2 introduces the deformation of director patterns and the conditions for metric compatibility. Sections 3 and 4 analyze the kinematics of three-fold and four-fold vertices, respectively. Section 5 presents two design strategies for large-scale director fields and their actuated configurations. Finally, Section 6 concludes the paper and discusses future research directions.
Notation. $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{3}\}$ is the standard right-handed orthonormal basis for $\mathbb{R}^{3}$. ${\bf R}_{{\bf e}_{3}}(\theta)$ is a rotation tensor with rotation axis ${\bf e}_{3}$ and rotation angle $\theta$. The sector angle is denoted as $\alpha_{i}$ on the reference domain and $\tilde{\alpha}_{i}$ on the actuated domain respectively.
2 Deformation of constant director fields and metric compatibility condition
We first consider the deformation induced by actuating a director pattern
${\bf n}({\bf x})=\cos\theta({\bf x}){\bf e}_{1}+\sin\theta({\bf x}){\bf e}_{2},~{}{\bf x}\in\mathbb{R}^{2}$ (1)
in the ${\bf e}_{1},{\bf e}_{2}$ plane, where ${\bf n}({\bf x})$ is a unit vector. Upon actuation, the pattern contracts along the director ${\bf n}$ by a factor $\lambda_{\parallel}$ and elongates along the perpendicular direction ${\bf n}_{\perp}$ by a factor $\lambda_{\perp}$. Here ${\bf n}_{\perp}$ is also called the orthogonal dual of ${\bf n}$ satisfying ${\bf n}_{\perp}\cdot{\bf n}=0$. Specifically, in liquid crystal elastomers/glasses, the contraction factor $\lambda_{\parallel}=\lambda$ is $\sim 0.7$ and the elongation factor $\lambda_{\perp}=\lambda^{-\nu}$, with the opto-thermal Poisson’s ratio $\nu=1/2$ for incompressible elastomers and as high as $\nu=2$ for glasses. Encoding the contraction and elongation, one can write the deformation gradient of the actuation as
${\bf U}_{{\bf n}}=\lambda_{\parallel}{\bf n}\otimes{\bf n}+\lambda_{\perp}{\bf n}_{\perp}\otimes{\bf n}_{\perp}+{\bf e}_{3}\otimes{\bf e}_{3},$ (2)
which implies ${\bf U}_{{\bf n}}{\bf n}=\lambda_{\parallel}{\bf n}$ and ${\bf U}_{{\bf n}}{\bf n}_{\perp}=\lambda_{\perp}{\bf n}_{\perp}$. Notice that ${\bf U}_{{\bf n}}$ is invariant under the transformation ${\bf n}\rightarrow-{\bf n}$, which means the director ${\bf n}$ is actually double-arrowed, and changing ${\bf n}$ to $-{\bf n}$ results in the same pattern and actuation.
Figure 2: (a) A disc with constant director is deformed into an ellipse. (b) A sector angle $\alpha$ is deformed to $\tilde{\alpha}$. (c) A metric compatible interface with tangent ${\bf t}$ (or ${\bf t}^{\prime}$) between two directors ${\bf n}_{1}$ and ${\bf n}_{2}$.
The deformation of a constant director pattern is shown in Fig. 2(a), where a unit disc is deformed into an ellipse with long axis $\lambda_{\perp}$ and short axis $\lambda_{\parallel}$ under ${\bf U}_{{\bf n}}$. Accordingly, the change of sector angles for a circular sector before and after actuation can be computed. The sector angle $\alpha\in(0,\pi)$ and two unit side vectors ${\bf t}_{1}$, ${\bf t}_{2}$ on the reference domain are illustrated in Fig. 2(b). Under the deforamtion gradient ${\bf U}_{{\bf n}}$, the reference and deformed sector angles $\alpha$ and $\tilde{\alpha}$ are given by
$\alpha=\alpha({\bf t}_{1},{\bf t}_{2})=\arccos({\bf t}_{1}\cdot{\bf t}_{2}),\quad\tilde{\alpha}=\tilde{\alpha}({\bf t}_{1},{\bf t}_{2},{\bf n})=\arccos\left(\frac{{\bf U}_{{\bf n}}{\bf t}_{1}\cdot{\bf U}_{{\bf n}}{\bf t}_{2}}{|{\bf U}_{{\bf n}}{\bf t}_{1}||{\bf U}_{{\bf n}}{\bf t}_{2}|}\right)$ (3)
respectively.
Next, we consider the metric-compatible interface with tangent ${\bf t}$ between the two constant director fields ${\bf n}_{1}$ and ${\bf n}_{2}$ satisfying ${\bf n}_{1}\nparallel{\bf n}_{2}$. The interface will serve as the crease of origami in the following sections. If we assume that the interfaces are compatible before and after actuation (the lengths are identical from the two sides), one can obtain the metric-compatibility condition
$|{\bf U}_{{\bf n}_{1}}{\bf t}|=|{\bf U}_{{\bf n}_{2}}{\bf t}|\Leftrightarrow({\bf n}_{1}\cdot{\bf t})^{2}=({\bf n}_{2}\cdot{\bf t})^{2},$ (4)
which yields the unit tangent ${\bf t}$ as
${\bf t}=\frac{\sigma{\bf n}_{1}+\tilde{\sigma}{\bf n}_{2}}{|\sigma{\bf n}_{1}+\tilde{\sigma}{\bf n}_{2}|},\quad\sigma,\tilde{\sigma}\in\{\pm 1\}.$ (5)
Figure 2(c) shows an example with ${\bf t}=({\bf n}_{1}+{\bf n}_{2})/|{\bf n}_{1}+{\bf n}_{2}|$ on the reference domain. By satisfying the condition (5), the deformed interfaces can be joined together with tangent $\tilde{{\bf t}}$ by appropriate rigid motions. It should be noted that, given the two director fields, the metric-compatible interface is either a bisector or a perpendicular dual to the bisector (${\bf t}$ and ${\bf t}^{\prime}$ in Fig. 2(c)).
3 Three-fold intersection
In this section, we study the kinematics of three-fold intersections with patterned piecewise constant director fields satisfying the metric compatibility condition at each interface. Our approach is shown in Fig. 3(a). The reference pattern is metric compatible at each interface. To compute the actuated configuration, the first step is to cut the reference pattern along the interface ${\bf t}_{1}$ and then deform it by the spontaneous deformation gradient ${\bf U}_{{\bf n}_{i}}$ at each domain. This will form an intermediate configuration in 2D, with a deficit or a surplus. The second step is to deform the intermediate configuration isometrically and stitch the boundaries together to form a non-Euclidean origami with Gaussian curvature at the tip. We use $\alpha_{i}$ and $\tilde{\alpha}_{i}$ to denote the sector angles before and after actuation, respectively.
Figure 3: (a) A three-fold non-Euclidean origami constructed through cutting, deforming by ${\bf U}_{{\bf n}}$, and stitching. (b) A three-fold intersection with a deficit ($\tilde{\alpha}_{1}+\tilde{\alpha}_{2}+\tilde{\alpha}_{3}<2\pi$). (c) A three-fold intersection with a surplus ($\tilde{\alpha}_{1}+\tilde{\alpha}_{2}+\tilde{\alpha}_{3}>2\pi$).
3.1 Three-fold origami with angular deficit and surplus
We first employ the continuum mechanics approach to derive the kinematics from the intermediate state to the folded state for the three-fold origami, where the intermediate state has an angular deficit or surplus. Let $\tilde{{\bf x}}_{i}=\tilde{\delta}_{i}\tilde{{\bf t}}_{i}$, where $\tilde{{\bf t}}_{i}$ is the unit tangent of the crease, for $i=1,2,3,4$, in the intermediate state. Let the sector angle between $\tilde{{\bf t}}_{i}$ and $\tilde{{\bf t}}_{i+1}$ be $\tilde{\alpha}_{i}$, as shown in Fig. 3(b). Then the tangents are given by
$\tilde{{\bf t}}_{1}={\bf e}_{1},~{}\tilde{{\bf t}}_{2}={\bf R}_{{\bf e}_{3}}(\tilde{\alpha}_{1}){\bf e}_{1},~{}\tilde{{\bf t}}_{3}={\bf R}_{{\bf e}_{3}}(\tilde{\alpha}_{1}+\tilde{\alpha}_{2}){\bf e}_{1},~{}\tilde{{\bf t}}_{4}={\bf R}_{{\bf e}_{3}}(\tilde{\alpha}_{1}+\tilde{\alpha}_{2}+\tilde{\alpha}_{3}){\bf e}_{1}.$ (6)
The pattern with an angular deficit (surplus) satisfies $\tilde{\alpha}_{1}+\tilde{\alpha}_{2}+\tilde{\alpha}_{3}<2\pi$ ($\tilde{\alpha}_{1}+\tilde{\alpha}_{2}+\tilde{\alpha}_{3}>2\pi$).
By folding along the creases isometrically, the origami may form a continuous configuration in 3D by matching $\tilde{{\bf t}}_{4}$ and $\tilde{{\bf t}}_{1}$, as illustrated in Fig. 3. We employ the continuum mechanics approach to derive the folding angles to achieve the continuous configuration. The deformation gradients of the panels are ${\bf I}$, ${\bf R}_{2}(\gamma_{2})$, and ${\bf R}_{2}(\gamma_{2}){\bf R}_{3}(\gamma_{3})$ respectively, where ${\bf R}_{i}(.)=\cos(.)(\tilde{{\bf t}}_{i}\otimes\tilde{{\bf t}}_{i}+\tilde{{\bf t}}_{i}^{\perp}\otimes\tilde{{\bf t}}_{i}^{\perp})+\sin(.)(-\tilde{{\bf t}}_{i}\otimes\tilde{{\bf t}}_{i}^{\perp}+\tilde{{\bf t}}_{i}^{\perp}\otimes\tilde{{\bf t}}_{i})+{\bf e}_{3}\otimes{\bf e}_{3}$ is a rotation tensor about $\tilde{{\bf t}}_{i}$. Here $\tilde{{\bf t}}_{i}^{\perp}={\bf R}_{{\bf e}_{3}}(\pi/2)\tilde{{\bf t}}_{i}$. The continuity at $\tilde{{\bf t}}_{1}$ (and $\tilde{{\bf t}}_{4}$) requires the compatibility condition
${\bf R}_{2}(\gamma_{2}){\bf R}_{3}(\gamma_{3})\tilde{{\bf t}}_{4}=\tilde{{\bf t}}_{1}$ (7)
for both the deficit case and the surplus case. Satisfying the compatibility condition yields a folding deformation ${\bf y}(\tilde{{\bf x}})$ that maps the intermediate state to the folded state as
${\bf y}(\tilde{{\bf x}})=\left\{\begin{array}[]{ll}\tilde{{\bf x}},&\tilde{{\bf x}}\cdot{\bf e}_{3}=0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}_{2}^{\perp}<0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}_{1}^{\perp}\geq 0,\\ {\bf R}_{2}(\gamma_{2})\tilde{{\bf x}},&\tilde{{\bf x}}\cdot{\bf e}_{3}=0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}_{3}^{\perp}<0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}_{2}^{\perp}\geq 0,\\ {\bf R}_{2}(\gamma_{2}){\bf R}_{3}(\gamma_{3})\tilde{{\bf x}},&\tilde{{\bf x}}\cdot{\bf e}_{3}=0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}_{4}^{\perp}<0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}_{3}^{\perp}\geq 0.\end{array}\right.$ (8)
The remaining problem is to determine the folding angles $\gamma_{2}$ and $\gamma_{3}$, which are given by the following theorem. The theorem implies that there are two compatible deformed configurations at fixed folding angles $\gamma_{2}$ and $\gamma_{3}$.
Theorem 3.1.
Given $\tilde{\alpha}_{1},\tilde{\alpha}_{2},\tilde{\alpha}_{3}\in(0,2\pi)\setminus\{\pi\}$ as sector angles, the compatibility condition (7) holds if and only if
$\gamma_{2}$ $=$ $\sigma\arccos\left(\frac{-\cos\tilde{\alpha}_{3}+\cos\tilde{\alpha}_{1}\cos\tilde{\alpha}_{2}}{\sin\tilde{\alpha}_{1}\sin\tilde{\alpha}_{2}}\right),$
$\gamma_{3}$ $=$ $\sigma\operatorname{sign}\left(\frac{\sin\tilde{\alpha}_{1}}{\sin\tilde{\alpha}_{3}}\right)\arccos\left(\frac{-\cos\tilde{\alpha}_{1}+\cos\tilde{\alpha}_{2}\cos\tilde{\alpha}_{3}}{\sin\tilde{\alpha}_{2}\sin\tilde{\alpha}_{3}}\right),$ (9)
for $\tilde{\alpha}_{i}$ in the admissible shown in Fig. 4(a), where $\sigma\in{\pm 1}$ denotes the two branches.
Proof.
Projecting both sides of Eq. (7) on $\tilde{{\bf t}}_{2}$ yields
${\bf R}_{3}(\gamma_{3})\tilde{{\bf t}}_{4}\cdot\tilde{{\bf t}}_{2}=\tilde{{\bf t}}_{1}\cdot\tilde{{\bf t}}_{2}\Leftrightarrow\cos\gamma_{3}=\frac{-\cos\tilde{\alpha}_{1}+\cos\tilde{\alpha}_{2}\cos\tilde{\alpha}_{3}}{\sin\tilde{\alpha}_{2}\sin\tilde{\alpha}_{3}}.$ (10)
Similarly, dotting Eq. (7) with ${\bf R}_{2}^{\mathrm{T}}\tilde{{\bf t}}_{3}$ gives
$\cos\gamma_{2}=\frac{-\cos\tilde{\alpha}_{3}+\cos\tilde{\alpha}_{1}\cos\tilde{\alpha}_{2}}{\sin\tilde{\alpha}_{1}\sin\tilde{\alpha}_{2}}.$ (11)
To determine the sign of $\gamma_{2}$ and $\gamma_{3}$, we dot (7) with ${\bf e}_{3}$ and substitute $\cos\gamma_{i}$ to obtain
$\sin\gamma_{3}=\frac{\sin\tilde{\alpha}_{1}}{\sin\tilde{\alpha}_{3}}\sin\gamma_{2}.$ (12)
Thus the solution of $\gamma_{2}$ and $\gamma_{3}$ can be summarized as
$\gamma_{2}$ $=$ $\sigma\arccos\left(\frac{-\cos\tilde{\alpha}_{3}+\cos\tilde{\alpha}_{1}\cos\tilde{\alpha}_{2}}{\sin\tilde{\alpha}_{1}\sin\tilde{\alpha}_{2}}\right),$
$\gamma_{3}$ $=$ $\sigma\operatorname{sign}\left(\frac{\sin\tilde{\alpha}_{1}}{\sin\tilde{\alpha}_{3}}\right)\arccos\left(\frac{-\cos\tilde{\alpha}_{1}+\cos\tilde{\alpha}_{2}\cos\tilde{\alpha}_{3}}{\sin\tilde{\alpha}_{2}\sin\tilde{\alpha}_{3}}\right),$ (13)
where $\sigma\in\{\pm 1\}$ denotes the two branches. Since $\cos\gamma_{2},\cos\gamma_{3}\in[-1,1]$, we have the admissible domain (Fig. 4(a)) for $\tilde{\alpha}_{1,2,3}$ such that the absolute values of Eqs. (10) and (11) are less than or equal to 1. ∎
Figure 4: (a) The admissible domain of $\tilde{\alpha}_{1}$, $\tilde{\alpha}_{2}$ and $\tilde{\alpha}_{3}$. (b) Notation for a three-fold director field.
3.2 Connection to director fields
Now we consider the relationship between the reference director field and the final folded configuration. The final state can be determined directly from the initial director fields, eliminating the need to compute the intermediate state.
Let ${\bf n}_{i}={\bf R}_{{\bf e}_{3}}(\theta_{i}){\bf e}_{1}$, $i=1,2,3$ be the constant director on the $i-$th domain ${\cal S}_{i}$ in a three-fold field. The three domains are separated by three interfaces with unit tangents ${\bf t}_{i},i=1,2,3$, where ${\bf t}_{i}$ is the tangent of the interface between the domain ${\cal S}_{i}$ and ${\cal S}_{i-1}$ (here we define ${\cal S}_{0}={\cal S}_{3}$). Following the metric-compatible condition (4) and the tangent solution (5), the unit tangent ${\bf t}_{i}$ of the interface satisfies
${\bf t}_{i}=\frac{\sigma_{i}{\bf n}_{i}+\tilde{\sigma}_{i}{\bf n}_{i-1}}{|\sigma_{i}{\bf n}_{i}+\tilde{\sigma}_{i}{\bf n}_{i-1}|},$ (14)
where $\sigma_{i},\tilde{\sigma}_{i}\in\{\pm 1\}$. Equation (14) implies that there are four solutions of ${\bf t}_{2}$ for the prescribed ${\bf n}_{1}$ and ${\bf n}_{2}$. However, these tangents cannot be arbitrarily chosen from the four solutions when patterning a three-fold pattern--they need to obey the following topological condition
$\operatorname{sign}[({\bf t}_{1}\times{\bf t}_{2})\cdot{\bf e}_{3}]=\operatorname{sign}[({\bf t}_{2}\times{\bf t}_{3})\cdot{\bf e}_{3}]=\operatorname{sign}[({\bf t}_{3}\times{\bf t}_{1})\cdot{\bf e}_{3}]=+1$ (15)
to arrange ${\bf t}_{1},{\bf t}_{2},{\bf t}_{3}$ counterclockwise and ensure that the sector angles are less than $\pi$.
Denote the three vertices of the three-fold intersection by ${\bf x}_{i}=\delta_{i}{\bf t}_{i}$ (no summation), where $\delta_{i}>0$ is the length of the crease with tangent ${\bf t}_{i}$, $i=1,2,3$. Upon actuation, the panels remain flat and exhibit pure rotations about the crease. The three-fold intersection will pop up (or down) and form a dome-like structure if the sum of the actuated sector angles is less than $2\pi$ (Fig. 4(b)). We derive the non-isometric deformation from its reference state to the final deformed state. First, employing Eq. (3), one can obtain the deformed sector angles as
$\tilde{\alpha}_{i}=\arccos\left(\frac{{\bf U}_{{\bf n}_{i}}{\bf t}_{i}\cdot{\bf U}_{{\bf n}_{i}}{\bf t}_{i+1}}{|{\bf U}_{{\bf n}_{i}}{\bf t}_{i}||{\bf U}_{{\bf n}_{i}}{\bf t}_{i+1}|}\right),$ (16)
where $i=1,2,3$, ${\bf t}_{4}:={\bf t}_{1}$, and ${\bf t}_{i}$ is given by (14). We then substitute the deformed sector angles $\tilde{\alpha}_{i}$ into the folding angle function (9), to obtain the folding angles $\gamma_{i}$ in terms of the direction of the director $\theta_{i}$, as stated in the following theorem.
Figure 5: Director patterns consisting of two three-fold intersections with metric-compatible interfaces. The deformed configurations remain compatible when changing ${\bf n}_{4}$ but keeping ${\bf n}_{1},{\bf n}_{2},{\bf n}_{3}$ unchanged.
Theorem 3.2.
Assume that the three-fold intersection of piecewise constant director fields shown in Fig. 4(b) satisfies the topological condition (15) and the tangent condition (14). The directors are given by ${\bf n}_{i}={\bf R}_{{\bf e}_{3}}(\theta_{i}){\bf e}_{1}$, $i=1,2,3$. Then the folding angles $\gamma_{2},\gamma_{3}$ satisfying the compatibility condition (7) are given by
$\cos\gamma_{i}=(r^{2})^{-\hat{\sigma}_{1}\hat{\sigma}_{2}\hat{\sigma}_{3}}+\left(\hat{\sigma}_{i-1}\hat{\sigma}_{i+1}-\hat{\sigma}_{i-1}\hat{\sigma}_{i+1}(r^{2})^{-\hat{\sigma}_{1}\hat{\sigma}_{2}\hat{\sigma}_{3}}\right)\cos(\theta_{i}-\theta_{i-1}),$ (17)
where $\hat{\sigma}_{i}=\sigma_{i}\tilde{\sigma}_{i}$. This leads to two solutions for folding angles $\gamma_{2}$ and $\gamma_{3}$:
$\gamma_{i}=\pm\arccos\left[(r^{2})^{-\hat{\sigma}_{1}\hat{\sigma}_{2}\hat{\sigma}_{3}}+\left(\hat{\sigma}_{i-1}\hat{\sigma}_{i+1}-\hat{\sigma}_{i-1}\hat{\sigma}_{i+1}(r^{2})^{-\hat{\sigma}_{1}\hat{\sigma}_{2}\hat{\sigma}_{3}}\right)\cos(\theta_{i}-\theta_{i-1})\right],~{}i=2,3,$ (18)
where $\pm$ corresponds to the “pop-up” or the “pop-down” deformation respectively, and $r=\lambda_{\parallel}/\lambda_{\perp}$.
Proof.
We take $\cos\gamma_{2}$ as an example for the proof, and the other $\gamma_{i}$ can be obtained by cycling the indices. As shown in (11), we obtain the folding angle $\gamma_{2}$ as a function of the intermediate sector angles
$\cos\gamma_{2}=\frac{-\cos\tilde{\alpha}_{3}+\cos\tilde{\alpha}_{1}\cos\tilde{\alpha}_{2}}{\sin\tilde{\alpha}_{1}\sin\tilde{\alpha}_{2}},$ (19)
where the intermediate sector angles are given by (16). Specifically, we have
$\cos\tilde{\alpha}_{1}=\frac{{\bf U}_{{\bf n}_{1}}{\bf t}_{1}\cdot{\bf U}_{{\bf n}_{1}}{\bf t}_{2}}{|{\bf U}_{{\bf n}_{1}}{\bf t}_{1}||{\bf U}_{{\bf n}_{1}}{\bf t}_{2}|},~{}\cos\tilde{\alpha}_{2}=\frac{{\bf U}_{{\bf n}_{2}}{\bf t}_{2}\cdot{\bf U}_{{\bf n}_{2}}{\bf t}_{3}}{|{\bf U}_{{\bf n}_{2}}{\bf t}_{2}||{\bf U}_{{\bf n}_{2}}{\bf t}_{3}|},~{}\cos\tilde{\alpha}_{3}=\frac{{\bf U}_{{\bf n}_{3}}{\bf t}_{1}\cdot{\bf U}_{{\bf n}_{3}}{\bf t}_{3}}{|{\bf U}_{{\bf n}_{3}}{\bf t}_{1}||{\bf U}_{{\bf n}_{3}}{\bf t}_{3}|}.$ (20)
Notice that $|{\bf U}_{{\bf n}_{1}}{\bf t}_{2}|=|{\bf U}_{{\bf n}_{2}}{\bf t}_{2}|$, $|{\bf U}_{{\bf n}_{2}}{\bf t}_{3}|=|{\bf U}_{{\bf n}_{3}}{\bf t}_{3}|$, and $|{\bf U}_{{\bf n}_{1}}{\bf t}_{1}|=|{\bf U}_{{\bf n}_{3}}{\bf t}_{1}|$ by the metric compatibility. Substituting (20) into (19) yields
$\cos\gamma_{2}=\frac{({\bf U}_{{\bf n}_{1}}{\bf t}_{1}\cdot{\bf U}_{{\bf n}_{1}}{\bf t}_{2})({\bf U}_{{\bf n}_{2}}{\bf t}_{2}\cdot{\bf U}_{{\bf n}_{2}}{\bf t}_{3})-({\bf U}_{{\bf n}_{3}}{\bf t}_{1}\cdot{\bf U}_{{\bf n}_{3}}{\bf t}_{3})|{\bf U}_{{\bf n}_{2}}{\bf t}_{2}|^{2}}{\lambda_{\parallel}^{2}\lambda_{\perp}^{2}({\bf t}_{1}\times{\bf t}_{2})\cdot({\bf t}_{2}\times{\bf t}_{3})}.$ (21)
The denominator of (21) follows the equalities
$|{\bf U}_{{\bf n}_{1}}{\bf t}_{1}||{\bf U}_{{\bf n}_{1}}{\bf t}_{2}|\sin\tilde{\alpha}_{1}=\lambda_{\parallel}\lambda_{\perp}\sin\alpha_{1}=\lambda_{\parallel}\lambda_{\perp}({\bf t}_{1}\times{\bf t}_{2})\cdot{\bf e}_{3},$ (22)
$|{\bf U}_{{\bf n}_{2}}{\bf t}_{2}||{\bf U}_{{\bf n}_{2}}{\bf t}_{3}|\sin\tilde{\alpha}_{2}=\lambda_{\parallel}\lambda_{\perp}\sin\alpha_{2}=\lambda_{\parallel}\lambda_{\perp}({\bf t}_{2}\times{\bf t}_{3})\cdot{\bf e}_{3}.$
Substituting the expressions of ${\bf t}_{i}$ and ${\bf U}_{{\bf n}_{i}}$ into (21), in principal we have 64 different expressions for $\cos\gamma_{2}$, depending on the different choices of $\sigma_{i}$ and $\tilde{\sigma}_{i}$. However, we can reduce the number significantly. Notice that the set of choices $\{\{\hat{\sigma}_{1}{\bf t}_{1},\hat{\sigma}_{2}{\bf t}_{2},\hat{\sigma}_{3}{\bf t}_{3}\},\hat{\sigma}_{i}\in\{\pm 1\},i=1,2,3\}$ all give the same result for (21). Therefore, we only need to check 8 cases. Let
${\bf t}_{1}=\hat{\sigma}_{1}{\bf n}_{3}+{\bf n}_{1},~{}{\bf t}_{2}=\hat{\sigma}_{2}{\bf n}_{1}+{\bf n}_{2},~{}{\bf t}_{3}=\hat{\sigma}_{3}{\bf n}_{2}+{\bf n}_{3},$ (23)
for $\hat{\sigma}_{i}\in\{\pm 1\}$. The direct computation leads to 8 solutions for $\cos\gamma_{2}$, classified as
$\cos\gamma_{2}=\begin{cases}\frac{({\lambda_{\parallel}}^{2}-{\lambda_{\perp}}^{2})\cos({\theta_{2}}-{\theta_{1}})+{\lambda_{\perp}}^{2}}{{\lambda_{\parallel}}^{2}},\quad\text{if}~{}\{\hat{\sigma}_{1},\hat{\sigma}_{2},\hat{\sigma}_{3}\}=\{+1,+1,+1\}~{}\text{or}~{}\{-1,+1,-1\},\\ \frac{({\lambda_{\parallel}}^{2}-{\lambda_{\perp}}^{2})\cos({\theta_{2}}-{\theta_{1}})+{\lambda_{\parallel}}^{2}}{{\lambda_{\perp}}^{2}},\quad\text{if}~{}\{\hat{\sigma}_{1},\hat{\sigma}_{2},\hat{\sigma}_{3}\}=\{-1,+1,+1\}~{}\text{or}~{}\{+1,+1,-1\},\\ \frac{(-{\lambda_{\parallel}}^{2}+{\lambda_{\perp}}^{2})\cos({\theta_{2}}-{\theta_{1}})+{\lambda_{\parallel}}^{2}}{{\lambda_{\perp}}^{2}},\quad\text{if}~{}\{\hat{\sigma}_{1},\hat{\sigma}_{2},\hat{\sigma}_{3}\}=\{+1,-1,+1\}~{}\text{or}~{}\{-1,-1,-1\},\\ \frac{(-{\lambda_{\parallel}}^{2}+{\lambda_{\perp}}^{2})\cos({\theta_{2}}-{\theta_{1}})+{\lambda_{\perp}}^{2}}{{\lambda_{\parallel}}^{2}},\quad\text{if}~{}\{\hat{\sigma}_{1},\hat{\sigma}_{2},\hat{\sigma}_{3}\}=\{-1,-1,+1\}~{}\text{or}~{}\{+1,-1,-1\}.\end{cases}$ (24)
Summarizing all the above solutions, we have a unified expression
$\cos\gamma_{2}=(r^{2})^{-\hat{\sigma}_{1}\hat{\sigma}_{2}\hat{\sigma}_{3}}+\left(\hat{\sigma}_{1}\hat{\sigma}_{3}-\hat{\sigma}_{1}\hat{\sigma}_{3}(r^{2})^{-\hat{\sigma}_{1}\hat{\sigma}_{2}\hat{\sigma}_{3}}\right)\cos(\theta_{2}-\theta_{1}),$ (25)
where $r=\lambda_{\parallel}/\lambda_{\perp}$. For all the 64 cases, one can replace $\hat{\sigma}_{i}$ by $\sigma_{i}\tilde{\sigma}_{i}$, which leads to Eq. (17) for $i=2$. Cycling $i=1,2,3$ in (25) yields other folding angles and completes the proof. ∎
An interesting observation from Theorem 3.2 is that, fixing $\sigma_{i}$ and $\tilde{\sigma}_{i}$, the folding angle $\gamma_{i}$ depends only on the directors on its adjacent panels, i.e., ${\bf n}_{i}$ and ${\bf n}_{i-1}$ with the directional angles $\theta_{i}$ and $\theta_{i-1}$. This fact provides more freedom to design connected three-fold intersections. For example, in Fig. 5, the pattern consists of two three-fold intersections. We may keep ${\bf n}_{2}$ and ${\bf n}_{3}$ unchanged and vary the direction of ${\bf n}_{4}$. As a result of Theorem 3.2, the folding angle along the shared crease in the actuated configuration is independent of ${\bf n}_{4}$ and will remain unchanged. Thus, the actuated configuration containing the two three-fold nodes remains compatible. We will use the basic pattern in Fig. 5 to design more complex patterns in Section 5.
4 Four-fold intersection
In this section, we study the kinematics of a general non-Euclidean four-fold origami generated by piecewise constant director fields. Again, we deform the intermediate state isometrically to achieve a continuous folded state. The method we use to compute the folding angle functions is partially adapted from our previous work zou2024kinematics . Here, we use a new symmetry approach to make the folding angle functions rigorous and more convenient to use for computation. We also provide a more physically relevant understanding of the “branches” in Fig. 7.
4.1 Reference domain
Suppose the intermediate state is given in Figure 6(a), where $0<\tilde{\alpha}_{1},\tilde{\alpha}_{2},\tilde{\alpha}_{3},\tilde{\alpha}_{4}<\pi$ are the sector angles. The tangents of the creases are given by
$\tilde{{\bf t}}_{1}={\bf e}_{1},~{}\tilde{{\bf t}}_{2}={\bf R}_{{\bf e}_{3}}(\tilde{\alpha}_{1}){\bf e}_{1},~{}\tilde{{\bf t}}_{3}={\bf R}_{{\bf e}_{3}}(\tilde{\alpha}_{1}+\tilde{\alpha}_{2}){\bf e}_{1},~{}\tilde{{\bf t}}_{4}={\bf R}_{{\bf e}_{3}}(\tilde{\alpha}_{1}+\tilde{\alpha}_{2}+\tilde{\alpha}_{3}){\bf e}_{1},~{}\tilde{{\bf t}}_{5}={\bf R}_{{\bf e}_{3}}(\tilde{\alpha}_{1}+\tilde{\alpha}_{2}+\tilde{\alpha}_{3}+\tilde{\alpha}_{4}){\bf e}_{1}.$ (26)
In general, the sum of the four sector angles $\sum_{i=1}^{4}\alpha_{i}$ is not $2\pi$, resulting in nonzero Gaussian curvature at the tip in the actuated configuration. Also, we have $\tilde{{\bf t}}_{5}\neq\tilde{{\bf t}}_{1}$ and a deficit (or surplus) may exist in the reference domain. Alternatively, a second version of the intermediate domain can be obtained with a cut between $\tilde{{\bf t}}_{4}$ and $\tilde{{\bf t}}_{55}$ where $\tilde{{\bf t}}_{55}=\cos\tilde{\alpha}_{4}{\bf e}_{1}-\sin\tilde{\alpha}_{4}{\bf e}_{2}$, as shown in Figure 6(a). Combining the equivalent two domains is convenient for computing the folding angles in the next section.
Figure 6: (a) Two equivalent intermediate configurations. (b) Two folding branches denoted by $\sigma=+1$ and $\sigma=-1$.
[…]
${\bf y}(\tilde{{\bf x}})=\left\{\begin{array}[]{ll}\tilde{{\bf x}},&\tilde{{\bf x}}\cdot{\bf e}_{3}=0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}_{2}^{\perp}<0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}_{1}^{\perp}\geq 0,\\ {\bf R}_{2}(\gamma_{2})\tilde{{\bf x}},&\tilde{{\bf x}}\cdot{\bf e}_{3}=0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}_{3}^{\perp}<0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}_{2}^{\perp}\geq 0,\\ {\bf R}_{2}(\gamma_{2}){\bf R}_{3}(\gamma_{3})\tilde{{\bf x}},&\tilde{{\bf x}}\cdot{\bf e}_{3}=0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}_{4}^{\perp}<0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}_{3}^{\perp}\geq 0,\\ {\bf R}_{2}(\gamma_{2}){\bf R}_{3}(\gamma_{3}){\bf R}_{4}(\gamma_{4})\tilde{{\bf x}},&\tilde{{\bf x}}\cdot{\bf e}_{3}=0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}_{5}^{\perp}<0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}_{4}^{\perp}\geq 0.\end{array}\right.$ (27)
The deformation is continuous except at the $\tilde{{\bf t}}_{1}\tilde{{\bf t}}_{5}$ boundary. Thus, a compatibility condition is needed to ensure continuity, which leads to the solutions of the folding angles in the next section. Using the second version of the intermediate domain (Fig. 6(a) bottom), the deformation in the fourth sector domain $\{\tilde{{\bf x}}\cdot{\bf e}_{3}=0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}^{\perp}_{1}<0,\ \tilde{{\bf x}}\cdot\tilde{{\bf t}}^{\perp}_{55}\geq 0\}$ can be written as
${\bf y}(\tilde{{\bf x}})={\bf R}_{1}(-\gamma_{1})\tilde{{\bf x}},$ (28)
which includes the folding angle $\gamma_{1}$.
4.2 Compatibility condition and folding angle functions
To ensure the continuity at the $\tilde{{\bf t}}_{1}\tilde{{\bf t}}_{5}$ boundary, the deformation gradients need to satisfy the compatibility condition:
${\bf R}_{2}(\gamma_{2}){\bf R}_{3}(\gamma_{3}){\bf R}_{4}(\gamma_{4})\tilde{{\bf t}}_{5}=\tilde{{\bf t}}_{1}.$ (29)
The folding angles are restricted in the interval $-\pi\leq\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4}\leq\pi$. In the rigidly and flat-foldable case, the four folding angles reach ±$\pm\pi$ simultaneously, with a reference free folding angle ranging from $-\pi$ to $\pi$. In the non-Euclidean case, some folding angles will reach ±$\pm\pi$ first, making the domain of the reference folding angle not a full interval of $[-\pi,\pi]$.
4.2.1 Calculations of $\gamma_{2}$, $\gamma_{3}$ and $\gamma_{4}$
To solve the compatibility equation, we first multiply (29) by ${\bf R}_{2}^{T}(\gamma_{2})$ to have
${\bf R}_{3}(\gamma_{3}){\bf R}_{4}(\gamma_{4})\tilde{{\bf t}}_{5}={\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1},$ (30)
and then dot (30) with $\tilde{{\bf t}}_{3}$ to have a scalar equation:
${\bf R}_{4}(\gamma_{4})\tilde{{\bf t}}_{5}\cdot\tilde{{\bf t}}_{3}-{\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}\cdot\tilde{{\bf t}}_{3}=0$ (31)
(recalling ${\bf R}_{3}(\gamma_{3})\tilde{{\bf t}}_{3}=\tilde{{\bf t}}_{3}$). This explicitly gives
$\cos(\gamma_{4})$ $=$ $\frac{\cos\tilde{\alpha}_{3}\cos\tilde{\alpha}_{4}-\cos\tilde{\alpha}_{1}\cos\tilde{\alpha}_{2}+\sin\tilde{\alpha}_{1}\sin\tilde{\alpha}_{2}\cos{\gamma}_{2}}{\sin\tilde{\alpha}_{3}\sin\tilde{\alpha}_{4}}$ (32)
::=:= $F_{4}(\gamma_{2};\tilde{\boldsymbol{\alpha}}),$
where we define $\tilde{\boldsymbol{\alpha}}:=(\tilde{\alpha}_{1},\tilde{\alpha}_{2},\tilde{\alpha}_{3},\tilde{\alpha}_{4})$. It is convenient to choose $\gamma_{2}$ as the reference folding angle, say, $\gamma_{2}=\theta$. Then, $\gamma_{4}$ as a function of $\theta$ is given in the condensed form
$\gamma_{4}=\bar{\gamma}_{4}^{\sigma}(\theta):=\sigma\arccos F_{4}(\theta;\tilde{\boldsymbol{\alpha}}).$ (33)
Here $\sigma\in\{\pm\}$ denotes the two branches of solution as shown in Fig. 6(b). Since the function $\bar{\gamma}_{4}^{\sigma}(\theta)$ is even and monotonic in $[-\pi,0]$ and $[0,\pi]$, we may characterize the admissible domain of $\theta$ by ${\cal I}^{-}\cup{\cal I}^{+}:=[-\theta_{\max},-\theta_{\min}]\cup[\theta_{\min},\theta_{\max}]$ in the following cases:
  1. 1.
    $F_{4}(0;\tilde{\boldsymbol{\alpha}})\in[-1,1]$. For this case, $\theta_{\min}=0$, and $\theta_{\max}$ is given by
    $\theta_{\max}=\begin{cases}\pi,&\text{if}~{}F_{4}(\pi;\tilde{\boldsymbol{\alpha}})\geq-1\\ \arccos\left[\frac{\cos\tilde{\alpha}_{1}\cos\tilde{\alpha}_{2}-\cos(\tilde{\alpha}_{3}-\tilde{\alpha}_{4})}{\sin\tilde{\alpha}_{1}\sin\tilde{\alpha}_{2}}\right],&\text{if}~{}F_{4}(\pi;\tilde{\boldsymbol{\alpha}})<-1\end{cases}.$ (34)
  2. 2.
    $F_{4}(0;\tilde{\boldsymbol{\alpha}})>1$ and $F_{4}(\pi;\tilde{\boldsymbol{\alpha}})\leq 1$. For this case, by solving $F_{4}(\theta_{\min};\tilde{\boldsymbol{\alpha}})=1$ and $F_{4}(\theta_{\max};\tilde{\boldsymbol{\alpha}})=-1$, we have
    $\theta_{\min}=\arccos\left[\frac{\cos\tilde{\alpha}_{1}\cos\tilde{\alpha}_{2}-\cos(\tilde{\alpha}_{3}+\tilde{\alpha}_{4})}{\sin\tilde{\alpha}_{1}\sin\tilde{\alpha}_{2}}\right],$ (35)
    $\theta_{\max}=\begin{cases}\pi,\quad&\text{if}~{}F_{4}(\pi;\tilde{\boldsymbol{\alpha}})\geq-1\\ \arccos\left[\frac{\cos\tilde{\alpha}_{1}\cos-\tilde{\alpha}_{2}-\cos(\tilde{\alpha}_{3}-\tilde{\alpha}_{4})}{\sin\tilde{\alpha}_{1}\sin\tilde{\alpha}_{2}}\right],\quad&\text{if}~{}F_{4}(\pi;\tilde{\boldsymbol{\alpha}})<-1\end{cases}.$ (36)
  3. 3.
    Otherwise, if $F_{4}(0;\tilde{\boldsymbol{\alpha}})$ and $F_{4}(\pi;\tilde{\boldsymbol{\alpha}})$ are beyond the above range, the admissible set of $\theta$ is empty.
Next, we dot (29) with ${\bf R}_{2}^{T}(\gamma_{2})$ and then project onto the plane perpendicular to $\tilde{{\bf t}}_{3}$ to have
${\bf R}_{3}(\gamma_{3})(({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{4}(\bar{\gamma}_{4}^{\sigma}(\gamma_{2}))\tilde{{\bf t}}_{5})=({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}.$ (37)
The norms of the projected vectors are the same:
$|({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3})({\bf R}_{3}(\gamma_{3}){\bf R}_{4}(\bar{\gamma}_{4}^{\sigma}(\gamma_{2}))\tilde{{\bf t}}_{5}|=|({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}|,$ (38)
which is trivially true. Thus, to solve the projected equation (37), we consider the non-degenerate case under $|({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}|\neq 0$. For (37) to hold, $\gamma_{3}=\bar{\gamma}_{3}^{\sigma}(\gamma_{2})$ must satisfy
$\cos(\bar{\gamma}_{3}^{\sigma}(\gamma_{2}))$ $=$ $\frac{({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{4}(\bar{\gamma}_{4}^{\sigma}(\gamma_{2}))\tilde{{\bf t}}_{5}\cdot({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}}{|({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{4}(\bar{\gamma}_{4}^{\sigma}(\gamma_{2}))\tilde{{\bf t}}_{5}||({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}|}$ (39)
::=:= $F_{3c}^{\sigma}(\gamma_{2};\tilde{\boldsymbol{\alpha}}),$
$\sin(\bar{\gamma}_{3}^{\sigma}(\gamma_{2}))$ $=$ $\text{sign}\left[\tilde{{\bf t}}_{3}\cdot({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{4}(\bar{\gamma}_{4}^{\sigma}(\gamma_{2}))\tilde{{\bf t}}_{5}\times({\bf I}-\tilde{{\bf t}}_{3}\otimes\tilde{{\bf t}}_{3}){\bf R}_{2}^{T}(\gamma_{2})\tilde{{\bf t}}_{1}\right]\sqrt{1-\cos^{2}(\bar{\gamma}_{3}^{\sigma}(\gamma_{2}))}$ (40)
::=:= $F_{3s}^{\sigma}(\gamma_{2};\tilde{\boldsymbol{\alpha}}).$
Then the solution of $\gamma_{3}$ is given by
$\gamma_{3}=\bar{\gamma}_{3}^{\sigma}(\gamma_{2})=\text{sign}[F_{3s}^{\sigma}(\gamma_{2};\tilde{\boldsymbol{\alpha}})]\arccos[F_{3c}^{\sigma}(\gamma_{2};\tilde{\boldsymbol{\alpha}})].$ (41)
Figure 7: Two examples with the same set of sector angles. (a) The folding angle functions show discontinuities by choosing $\gamma_{2}$ as the reference folding angle. (b) The folding angle functions are continuous by cycling sector angles and then choosing $\gamma_{2}$ (which is $\gamma_{1}$ in (a)) as the reference folding angle.
4.2.2 Calculation of $\gamma_{1}$
To solve $\gamma_{1}$, we use the second intermediate domain in Fig. 6(a) such that there exists $\gamma_{1}=\bar{\gamma}_{1}^{\sigma}(\theta)$ making the folded configuration continuous across $\tilde{{\bf t}}_{4}$ and $\tilde{{\bf t}}_{55}$. A convenient way to obtain $\gamma_{1}$ is to use the cyclic symmetry of $\tilde{\alpha}_{1},\tilde{\alpha}_{2},\tilde{\alpha}_{3}$ and $\tilde{\alpha}_{4}$. Notice that $\gamma_{4}=\bar{\gamma}_{4}^{\sigma}(\gamma_{2})$ shows the ability to solve the folding angle ($\gamma_{4}$) in terms of its opposite folding angle ($\gamma_{2}$). Thus we may solve $\gamma_{1}$ in terms of its opposite folding angle $\gamma_{3}$, by simply cycling $\tilde{\alpha}_{1},\tilde{\alpha}_{2},\tilde{\alpha}_{3},\tilde{\alpha}_{4}$ and then using the the same function form $F_{4}$ in (32) as follows:
$\gamma_{1}=\hat{\sigma}\arccos F_{4}[(\bar{\gamma_{3}}^{\sigma}(\gamma_{2};\tilde{\boldsymbol{\alpha}});\hat{\boldsymbol{\alpha}}]:=\bar{\gamma}_{1}^{\hat{\sigma}}(\gamma_{2},\sigma),$ (42)
where $\hat{\boldsymbol{\alpha}}=(\tilde{\alpha}_{2},\tilde{\alpha}_{3},\tilde{\alpha}_{4},\tilde{\alpha}_{1})$. Substituting $\gamma_{3}$ as a function of $\gamma_{2}$, we can still use $\gamma_{2}$ as the reference folding angle in (42). To determine the folding branch $\hat{\sigma}$, we perform the following continuity check for $\tilde{{\bf t}}_{4}$ and $\tilde{{\bf t}}_{55}$:
$\hat{\sigma}(\theta)=\begin{cases}+1,~{}\text{if}~{}{\bf R}_{1}(-\arccos F_{4}(\bar{\gamma_{3}}^{\sigma}(\theta;\tilde{\boldsymbol{\alpha}});\hat{\boldsymbol{\alpha}}))\tilde{{\bf t}}_{55}={\bf R}_{2}(\theta){\bf R}_{3}(\bar{\gamma}_{3}^{\sigma}(\theta))\tilde{{\bf t}}_{4}\\ -1,~{}\text{if}~{}{\bf R}_{1}(\arccos F_{4}(\bar{\gamma_{3}}^{\sigma}(\theta;\tilde{\boldsymbol{\alpha}});\hat{\boldsymbol{\alpha}}))\tilde{{\bf t}}_{55}={\bf R}_{2}(\theta){\bf R}_{3}(\bar{\gamma}_{3}^{\sigma}(\theta))\tilde{{\bf t}}_{4}\end{cases}.$ (43)
Then the deformation of the fourth domain is accordingly ${\bf y}(\tilde{{\bf x}})={\bf R}_{1}(-\gamma_{1})\tilde{{\bf x}}$.
4.2.3 Admissible domain to avoid self-intersection
In our previous analysis, we choose $\gamma_{2}$ as the reference folding angle, and the folding angle functions ensure the continuity of the deformation ${\bf y}(\tilde{{\bf x}})$. However, this analysis allows unphysical self-intersections, and one panel may be folded to pass through its adjacent panels. Thus, we need to determine the admissible domain of the reference folding angle. The domain ${\cal I}^{+}\cup{\cal I}^{-}$ is obtained only from the existence of $\gamma_{4}$. Therefore, the admissible domain is a subset of ${\cal I}^{+}\cup{\cal I}^{-}$ such that no self-intersection occurs at any crease.
Notice that the folding angle jumps between $-\pi$ and $\pi$ when the panel intersects with its adjacent panel. Thus, we perform the following check
$|\gamma_{i}(\bar{\theta}^{+})-\gamma_{i}(\bar{\theta}^{-})|=2\pi$ (44)
to determine the reference folding angle $\bar{\theta}$ at which the folding angles $\gamma_{i}$ reaches ±$\pm\pi$. Let ${\cal J}^{+}\subset{\cal I}^{+}=[\theta_{\min},\theta_{\max}]$ be the set of points at which one of the folding angles jumps, say,
${\cal J}^{+}:=\{\bar{\theta}\in{\cal I}^{+}:|\gamma_{i}(\bar{\theta}^{+})-\gamma_{i}(\bar{\theta}^{-})|=2\pi,~{}i=1,2,3\text{ or }4\}.$ (45)
Define
$\bar{\theta}_{\max}:=\begin{cases}\inf{\cal J}^{+},\quad&\text{if}~{}{\cal J}^{+}\neq\emptyset\\ \theta_{\max},\quad&\text{if}~{}{\cal J}^{+}=\emptyset\end{cases}$ (46)
as the upper bound of the physically relevant reference folding angle (i.e., no self-intersection) on ${\cal I}^{+}$. Similarly, the lower bound $\bar{\theta}_{\min}$ on ${\cal I}^{-}=[-\theta_{\max},-\theta_{\min}]$ is defined as
$\bar{\theta}_{\min}:=\begin{cases}\sup{\cal J}^{-},\quad&\text{if}~{}{\cal J}^{-}\neq\emptyset\\ -\theta_{\max},\quad&\text{if}~{}{\cal J}^{-}=\emptyset\end{cases}$ (47)
where ${\cal J}^{-}$ is given by
${\cal J}^{-}:=\{\theta\in{\cal I}^{-}:|\gamma_{i}(\theta^{+})-\gamma_{i}(\theta^{-})|=2\pi,~{}i=1,2,3~{}\text{or}~{}4\}.$ (48)
Then the admissible domain of the reference folding angle is $\bar{{\cal I}}^{+}\cup\bar{{\cal I}}^{-}$, where
¯$\bar{{\cal I}}^{+}=[\theta_{\min},\bar{\theta}_{\max}],\quad\bar{{\cal I}}^{-}=[\bar{\theta}_{\min},-\theta_{\min}].$ (49)
In reality, the domain can be obtained numerically by simply performing the above check, resulting in the folding angle curves in Fig. 7 that do not exceed ±$\pm\pi$.
In summary, the folding angle functions that make the folded configuration continuous with no self-intersections are given by
$\gamma_{1}=\bar{\gamma}_{1}^{\tilde{\sigma}(\theta)}(\theta,\sigma),~{}\gamma_{2}=\theta,~{}\gamma_{3}=\bar{\gamma}_{3}^{\sigma}(\theta),~{}\gamma_{4}=\bar{\gamma}_{4}^{\sigma}(\theta),~{}\sigma\in\{\pm 1\},~{}\theta\in\bar{{\cal I}}^{-}\cup\bar{{\cal I}}^{+}.$ (50)
The Mathematica code for simulating the deformation of a general non-Euclidean four-fold origami is available upon request.
Remark 1: Symmetry relation. By checking the folding angle functions, we claim that if $(\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4})$ is on the branch $\sigma=+$, then $(-\gamma_{1},-\gamma_{2},-\gamma_{3},-\gamma_{4})$ is on the other branch $\sigma=-$, as shown in Fig. 7. These two configurations are mirror related.
Remark 2: Choices of different reference folding angles. It is noticed that, for the same crease pattern, choosing the appropriate reference folding angle may result in continuous folding angle functions (Fig. 7(b)) rather than discontinuous ones (Fig. 7(a)). Using the notations in Fig. 7(a), it is numerically verified that the folding angle functions are continuous in terms of $\gamma_{2}$ if the sector angles satisfy
$\begin{cases}\tilde{\alpha}_{1}+\tilde{\alpha}_{2}\leq\tilde{\alpha}_{3}+\tilde{\alpha}_{4}\\ \tilde{\alpha}_{1}+\tilde{\alpha}_{2}+\tilde{\alpha}_{3}+\tilde{\alpha}_{4}\leq 2\pi\end{cases}\quad\text{or}\quad\begin{cases}\tilde{\alpha}_{1}+\tilde{\alpha}_{2}\geq\tilde{\alpha}_{3}+\tilde{\alpha}_{4}\\ \tilde{\alpha}_{1}+\tilde{\alpha}_{2}+\tilde{\alpha}_{3}+\tilde{\alpha}_{4}\geq 2\pi\end{cases}.$ (51)
Remark 3: Connection to director fields. In the four-fold case, we do not have a direct relation between the director pattern on the reference domain and the folding angles on the folded configuration, as in the three-fold case. However, the numerical approach is similar. First, a topology condition is needed ($({\bf t}_{i}\times{\bf t}_{i+1})\cdot{\bf e}_{3}=+1$) for the initial director pattern. Then we can calculate the deformed sector angles by $\tilde{\alpha}_{i}=\arccos[({\bf U}_{{\bf n}_{i}}{\bf t}_{i}\cdot{\bf U}_{{\bf n}_{i}}{\bf t}_{i+1})/(|{\bf U}_{{\bf n}_{i}}{\bf t}_{i}||{\bf U}_{{\bf n}_{i}}{\bf t}_{i+1}|)]$. Substituting these deformed sector angles into the folding angle functions yields the kinematics of the four-fold origami.
5 Patterning the three-fold and four-fold intersections
It is important to pattern the three-fold and four-fold intersections into active metamaterials, relying on the kinematics of single nodes. The active metamaterials are designed to be flat before actuation, and will actuate into an effective plate/shell with concentrated Gaussian curvature at nodes after actuation. Here, we provide two systematic designs: quadrilateral director patterns consisting of four-fold nodes and director patterns consisting of both three-fold and four-fold nodes.
5.1 Quadrilateral director patterns
Figure 8: (a) The director on the $(3,3)$ panel (in the red frame) is determined by the other directors and the folding angles $\gamma_{3},\gamma_{4}$. (b) A marching algorithm: the director pattern within the dashed red frame is determined by the boundary directors and the reference folding angle, following (a).
The first director pattern we consider is a quadrilateral pattern consisting of four-fold intersections. The starting point is a $3\times 3$ pattern with an unknown director at the $(3,3)$ panel (see Fig. 8(a)). The other directors, creases and folding branches are prescribed. Suppose a reference folding angle $\theta$ is given at a crease of the $(1,1)$ panel. Due to the one-DoF feature, the folding angles $\gamma_{3},\gamma_{4}$ are determined on the $(3,3)$ panel, as shown in the red frame. Meanwhile, to satisfy the metric compatibility, the tangents ${\bf t}_{1}({\bf n}),{\bf t}_{2}({\bf n})$ and the sector angles on the $(3,3)$ panel are functions of the unknown director ${\bf n}$. We will show that ${\bf n}$ is uniquely determined.
Notice that there are two ways to compute the actuated sector angle $\tilde{\alpha}_{1}$ for the $(3,3)$ panel (see Fig. 8(a)). On the one hand, the sector angle ${\alpha}_{3}$ is prescribed. $\alpha_{2}({\bf n})$and ${\alpha}_{4}({\bf n})$ are functions of ${\bf n}$. Then we can compute the actuated sector angles $\tilde{\alpha}_{3}$, $\tilde{\alpha}_{2}({\bf n})$ and $\tilde{\alpha}_{4}({\bf n})$. Together with the folding angles $\gamma_{3}$ and $\gamma_{4}$, the directions of the actuated tangents $\tilde{{\bf t}}_{1}({\bf n})$ and $\tilde{{\bf t}}_{2}({\bf n})$ can be calculated. By direct calculation, the angle $\tilde{\alpha}_{1}$ between $\tilde{{\bf t}}_{1}({\bf n})$ and $\tilde{{\bf t}}_{2}({\bf n})$ is
$\alpha_{1}^{\star}({\alpha}_{2}({\bf n}),{\alpha}_{3},{\alpha}_{4}({\bf n}),\gamma_{3},\gamma_{4})=\arccos[\cos\tilde{\alpha}_{2}({\bf n})\left(\cos\tilde{\alpha}_{3}\cos\tilde{\alpha}_{4}({\bf n})-\cos\gamma_{4}\sin\tilde{\alpha}_{3}\sin\tilde{\alpha}_{4}({\bf n})\right)$
$-\sin\tilde{\alpha}_{2}({\bf n})(\cos\gamma_{3}\sin\tilde{\alpha}_{3}\cos\tilde{\alpha}_{4}({\bf n})+\cos\gamma_{3}\cos\gamma_{4}\cos\tilde{\alpha}_{3}\sin\tilde{\alpha}_{4}({\bf n})-\sin\tilde{\alpha}_{4}({\bf n})\sin\gamma_{3}\sin\gamma_{4})],$ (52)
where $\tilde{\alpha}_{i}$ is the actuated sector angle corresponding to $\alpha_{i}$. On the other hand, the actuated sector angle $\tilde{\alpha}_{1}$ can be directly calculated from the $(3,3)$ panel, given by
$\alpha_{def}({{\bf t}}_{1}({\bf n}),{{\bf t}}_{2}({\bf n}),{\bf n})=\arccos\left[\frac{{\bf U}_{{\bf n}}{\bf t}_{1}({\bf n})\cdot{\bf U}_{{\bf n}}{\bf t}_{2}({\bf n})}{|{\bf U}_{{\bf n}}{\bf t}_{1}({\bf n})||{\bf U}_{{\bf n}}{\bf t}_{2}({\bf n})|}\right].$ (53)
Solving the equation
$\alpha_{1}^{\star}({\alpha}_{2}({\bf n}),{\alpha}_{3},{\alpha}_{4}({\bf n}),\gamma_{3},\gamma_{4})=\alpha_{def}({{\bf t}}_{1}({\bf n}),{{\bf t}}_{2}({\bf n}),{\bf n})$ (54)
yields a (generically unique) solution for ${\bf n}$, which can be used to develop a marching algorithm for the entire pattern.
To design a quadrilateral pattern, we prescribe the directors and the corresponding tangents at the left and bottom boundaries (panels with indices $(i,1),(i,2),(1,j),(2,j)$, i.e. the domain outside the red dashed lines in Fig. 8(b)). By entering a reference folding angle $\theta$ and applying Eq. (54), the pattern within the red dashed domain is determined. Once all directors are determined, we fix the tangent branches and compute the tangents. The coordinates of the vertices are then computed by these tangents. The difficulty we might have is that the process may break the topology of a quadrilateral mesh. Thus, a more realistic way is to start with a symmetric pattern and then perturb the boundary directors. Figure 8(b) is an example of irregular quadrilateral patterns. We start with a symmetric pattern with directors along ${\bf R}_{{\bf e}_{3}}(\pi/4){\bf e}_{1}$ and ${\bf R}_{{\bf e}_{3}}(3\pi/4){\bf e}_{1}$. As shown in Fig. 8(b), the symmetric pattern consists of Type A nodes and Type B nodes, which will deform into pyramids and saddles, respectively. We then perturb the directors at the boundary vertices and compute the entire pattern using Eq. (54). During the marching process, the folding angle branches are fixed by $\sigma=+1$ in Eq. (50), meaning the opposite folding angles have the same sign. The resulting quadrilateral pattern is irregular in both the reference and actuated configurations, and the distribution of Gaussian curvature is non-symmetric. In the deformed configuration, we observe alternating positive and negative Gaussian curvature at the four-fold vertices.
5.2 Director patterns combining three-fold and four-fold intersections
Figure 9: (a) Notation of the unit cell. (b) Notation of the four-fold intersection in the pattern. (c) The designed pattern composed of three-fold and four-fold intersections. The pattern is compatible in both the reference and actuated configurations.
The second pattern we design is a combination of three-fold and four-fold vertices as shown in Fig. 9. The unit cell is composed of two connected three-fold intersections inherited from the pattern in Fig. 5, with the notation shown in Fig. 9(a). The directors on adjacent unit cells satisfy ${\bf n}_{3}(i,j)={\bf n}_{2}(i+1,j)$ and ${\bf n}_{4}(i,j)={\bf n}_{1}(i,j+1)$. Then the boundaries of unit cells in Fig. 9(c) (i.e. thin lines) are not folding lines, and the two adjacent panels on two unit cells form a single panel. The actual folding lines in Fig. 9(c) are thick lines. We then deform the reference pattern by popping the unit cells up or down, indicated by $\delta=+1$ or $\delta=-1$ respectively. It is noted that if the distribution of $\delta$ satisfies
$\begin{cases}\delta(i,j)=+1,\quad\text{}i+j\text{ odd}\\ \delta(i,j)=-1,\quad\text{}i+j\text{ even}\end{cases},$ (55)
the deformed pattern is compatible, in particular at the four-fold intersection. Figure 9(b) shows a generic four-fold intersection in the pattern. We will show that the deformed four-fold intersection is compatible if the folding angles are computed from the three-fold intersections. For an arbitrary four-fold intersection in the pattern (Fig. 9(b)), we assume the directors are given by ${\bf n}_{i}={\bf R}_{{\bf e}_{3}}(\theta_{i}){\bf e}_{1},0<\theta_{1}<\theta_{2}<\theta_{3}<\theta_{4}<2\pi$, which satisfy the topological condition of creases. We compute the folding angle $\gamma_{i}$ using Eq. (17) from the three-fold kinematics
$\gamma_{i}=(-1)^{i}\arccos\left[r^{2}+(-1+r^{2})\cos(\theta_{i}-\theta_{i-1})\right],~{}i=1,2,3,4.$ (56)
Here we have already substituted the tangent branch $\hat{\sigma}_{i}$ for this particular pattern and $(-1)^{i}$ accounts for the pop-up or pop-down of the three-fold nodes. That is, the folding angles are inherited from the three-fold intersections, and the folds of the actuated four-fold nodes have alternating mountains and valleys folding lines. Then we perform the numerical validation for the compatibility equation at the four-fold vertex:
$error=\max_{0<\theta_{1}<\theta_{2}<\theta_{3}<\theta_{4}<2\pi}|{\bf R}_{2}(\gamma_{2}){\bf R}_{3}(\gamma_{3}){\bf R}_{4}(\gamma_{4})\tilde{{\bf t}}_{5}-\tilde{{\bf t}}_{1}|$ (57)
for any $\theta_{i}$ satisfying the topological condition ${\bf t}_{i}\times{\bf t}_{i+1}\cdot{\bf e}_{3}=1$, where ${\bf R}_{i}$ is a rotation tensor with rotation axis $\tilde{{\bf t}}_{i}$ for the intermediate configuration. The numerical result given by the global minimization NMaximize in Mathematica shows that $error=0$ in the domain $0<\theta_{1}<\theta_{2}<\theta_{3}<\theta_{4}<2\pi$, indicating that the four-fold vertex is compatible in our setting. Since compatibility holds for arbitrary four-fold nodes, we can successively stack different unit cells along the i- and j-directions to construct both the reference pattern and the actuated configuration, as shown in Fig. 9(c). In the actuated state, the configuration exhibits positive Gaussian curvature at the three-fold vertices and negative Gaussian curvature at the four-fold vertices. Notably, the design space for the director fields is quite large; the primary constraint is the preservation of the pattern’s topological structure.
6 Discussion
In this work, we have presented a geometric framework for designing complex active origami patterns, where each panel is equipped with a prescribed metric change. This metric change, induced by local actuation, leads to nontrivial Gaussian curvature at the vertices, rendering the origami non-Euclidean. Such systems can be physically realized using two-dimensional liquid crystal elastomer (LCE) sheets, where a patterned director field encodes the spatially varying metric change. By assuming that each panel remains flat and rigid after actuation, and by neglecting crease energy, we reduce the design challenge to a purely geometric problem: how to construct a piecewise constant director field that ensures compatibility in both the reference and actuated configurations.
To tackle this problem, we first analyze the fundamental condition of metric compatibility between adjacent director fields. We then develop a rigorous continuum mechanics framework to study the kinematics of isolated three-fold and four-fold vertices. Building on this vertex-level understanding, we construct two classes of large-scale director patterns that are compatible both before and after actuation. In other words, the directors satisfy metric compatibility at each crease, and the resulting global deformation is geometrically compatible. Although the general compatibility problem remains open, our two proposed designs demonstrate broad and flexible design spaces. The first is a quadrilateral director pattern with free directors on the two boundaries. Using a marching algorithm and a rigidity theorem, the director information propagates to interior nodes, producing a compatible actuated configuration characterized by alternating positive and negative Gaussian curvature at the vertices. The second design combines three-fold and four-fold intersections, leveraging a folding angle theorem that shows the folding angle at a three-fold node depends solely on its adjacent directors. In this configuration, positive Gaussian curvature arises at the three-fold vertices and negative curvature at the four-fold vertices after actuation.
Looking ahead, several promising directions merit further exploration. On the mechanics side, one could investigate the effective behavior of these patterns under external loads. For instance, drawing inspiration from effective plate theories developed for origami based on bar-and-hinge models XU2024105832 ; xu2025modelingcomputationeffectiveelastic ; filipov2017bar , it is natural to ask whether an analogous plate theory for non-Euclidean origami can be formulated, and what the embedded Gaussian curvature would add to the system. On the application side, our designs could serve as building blocks for energy-absorbing layers, deployable structures, or load-bearing elements, where the induced Gaussian curvature provides additional strength and stiffness, as seen in recent studies of non-Euclidean curved-fold actuators duffy2021shape ; feng2022interfacial ; feng2024geometry ; zhai2020situ . Finally, from a geometric perspective, the complete solution of the general compatibility problem—determining whether a given compatible piecewise-constant director field leads to compatible actuated configurations—remains open. Addressing this challenge may require an augmented non-isometric, piecewise-constant analog of the Nash embedding theorem nash1954c or a non-isometric extension of the crumpled paper theory conti2008confining .
Acknowledgments
This work is based on discussions with Prof. Paul Plucinsky at USC, while we were working together at the University of Minnesota. F.F. acknowledges the financial support from the National Natural Science Foundation of China (Grant No. 12472061).